Molding Fabrication Total Estimated total machine-hours used Estimated total fixed manufacturing overhead Estimated variable manufacturing overhead per machine-hour 1,500 2,500 $12,500 $ 2.40 4,000 $16,500 $29,000 3.20 Job Q $13,000 $11,500 Job P $23,000 $29,000 Direct materials Direct labor cost Actual machine-hours used: Molding Fabrication 2,700 1,600 4,300 1,800 006 I 3,700 Total Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.
Required: For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments. 14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establis selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (D not round intermediate calculations. Round your final answer to nearest whole dollar.) Job Q Job P Total price for the job Selling price per unit
Answer #1

Working notes:

  • Total cost = Direct material + Direct labor + Variable manufacturing overhead + Fixed manufacturing overhead
  • Total price = Total cost x 1.8
  • Selling price = Total price / Number of units

(1) Job P

Total cost ($) = 23,000 + 29,000 + (2,700 x 2.4) + (1,600 x 3.2) + 12,500 = 64,500 + 6,480 + 5,120 = 76,100

Total price ($) = 76,100 x 1.8 = 136,980

Selling price ($) = 136,980 / 20 = 6,849

(1) Job Q

Total cost ($) = 13,000 + 11,500 + (1,800 x 2.4) + (1,900 + 3.2) + 16,500 = 41,000 + 4,320 + 6,080 = 51,400

Total price ($) = 51,400 x 1.8 = 92,520

Selling price ($) = 92,520 / 30 = 3,084